OpenStudy (anonymous):

Write a polynomial function f of least degree that has rational coefficients, a leading coefficient of 1, and the given zeros. 1) i, 2 - √3 2) 1, 4, 1 + √2

OpenStudy (anonymous):

@Mertsj Please Help!

OpenStudy (mertsj):

If i is a root, -i is also a root. If 2-sqrt3 is a root then 2+sqrt3 is also a root.

OpenStudy (mertsj):

So the equation is: \[(x-i)(x+i)(x-(2-\sqrt{3}))(x-(2+\sqrt{3}))=0\]

OpenStudy (mertsj):

Now need to multiply out and simplify.

OpenStudy (anonymous):

How do you multiply this??

OpenStudy (mertsj):

What is (x-i)(x+i)?

OpenStudy (anonymous):

Isn't both -i and +i

OpenStudy (mertsj):